# Effectively Leveraging Momentum Terms in Stochastic Line Search Frameworks for Fast Optimization of Finite-Sum Problems

### v3

@@ 1 Introduction
State-of-the-art methods to tackle problems of this form are primarily based on stochastic gradient descent (SGD) [robbins1951] and its accelerated variants [polyak1964, nesterov1983]. The use of stochastic gradient methods allows to leverage data redundancy to approach good solutions while carrying out low-cost iterations. A thorough discussion on stochastic gradient methods and their application in data science is available in [bottou2018].
@@
Adaptive SGD approaches [duchi2011, tieleman2012, zeiler2012, kingma2015] have established through the years as solid choices in the considered scenario. In particular, Adam algorithm [kingma2015] is now largely considered the preferred optimizer in deep learning settings. In an alternative stream of research, variance-reduced approaches [schmidt2017, rie2013, gower2023] have been studied and shown to possess faster convergence rates than the base algorithm and in line with full-batch gradient descent. However, in practice, these methods require significant computational resources, either needing periodic full gradient computations or high memory usage; for this reason, variance reduction methods are often found to be impractical to use in real-world scenarios with deep architectures and very large datasets. Moreover, the assumptions made to obtain the nice convergence results have been shown not to be satisfied in deep learning scenarios. In fact, no improvement in the convergence rate [defazio2019] w.r.t. SGD is obtainable under more realistic assumptions.
@@
In more recent years, the focus turned to the analysis of stochastic gradient methods under a specific assumption that is reasonable with modern deep learning models: the interpolation regime. In this scenario, the convergence rate of SGD was finally shown to match that of full-batch methods [ma2018, vaswani2020], giving some justification to the observed behavior of these methods. Moreover, under interpolation assumptions it was possible to design well-grounded strategies to choose the step size in an adaptive manner [mutschler2020, loizou2021].
@@
Now, one of the most promising developments concerns the use of stochastic line searches. In particular, the classical Armijo rule was adapted to be successfully used in the incremental scenario [vaswani2019]. More recently, a nonmonotone version of the stochastic Armijo line search was proposed [galli2023]. With these techniques, not only the convergence rate of standard gradient descent is shown to be obtainable, but very encouraging numerical results were also observed.
@@
As widely recognized in the machine learning literature, the addition of momentum terms in the update rule has been known to be beneficial within SGD approaches [sutskever2013importance, sebbouh2021almost, tseng1998incremental, jelassi2022towards, gitman2019understanding]; typically, this is attributed to the stabilizing effect on the direction and the speed-up in low curvature regions. It is still not clear, however, how this kind of additional term, which modifies the search direction, could be suitably handled in line search based frameworks. A heuristic choice for the step size in presence of momentum terms under interpolation assumptions has been proposed by [wang2023], but no decrease condition is coupled with this strategy. On the other hand, a backtracking strategy has been recently suggested, ensuring the resulting direction is of descent for the considered mini-batch [fan2023]; for this latter approach, however, we will outline massive shortcomings both on the theoretical and on the computational sides.
@@
In this manuscript, we therefore deal with the above challenge. Firstly, we shed light on an intrinsic issue related to the momentum direction in the incremental regime; in order to overcome it, at least partially, we propose to exploit the concept of mini-batch persistency [fischetti2018], which we also observe being beneficial on its own in certain settings.
@@
We can then focus on a suitable rule to define the parameter $\beta$ associated with the momentum term in the direction update. The proposed method exploits the strong connection of momentum methods and classical nonlinear conjugate gradient algorithms [grippo2023introduction, Ch. 12]: in both cases, the search direction is a linear combination of the current gradients and the previous direction, and the border between the two families of approaches is thin. The update rule is combined with safeguarding strategies based on restarting [chan2022nonlinear] or subspace optimization [lapucci2024globally], that allow to guarantee crucial assumptions for the soundness of the resulting algorithm [lapucci2024convergenceconditionsstochasticline].
@@ 2 Preliminaries
The update rule is then given by $x^{k+1}=x^{k}+\alpha_{k}d_{k}$, where the step size $\alpha_{k}$ is often referred to as learning rate and $d_{k}$ is the search direction. A deep presentation of stochastic gradient methods for machine learning tasks is available in [bottou2018]. Under regularity assumptions, and using a suitable decreasing sequence of steps $\{\alpha_{k}\}$, the iterative scheme allows to guarantee convergence to stationarity in expectation. The associated worst-case complexity bound is $\mathcal{O}(\frac{1}{\epsilon^{4}})$ in the nonconvex setting [ghadimi2013], which becomes $\mathcal{O}(\frac{1}{\epsilon^{2}})$ and $\mathcal{O}(\frac{1}{\epsilon})$ under convexity and strong convexity assumptions, respectively [bottou2018]. These bounds are worse than those obtainable by full-batch methods, which should be thus superior. Yet, the empirical evidence collected training deep machine learning models tells another story.
Recently, some explanation to this gap between theory and numerical observations was found: modern machine learning models are usually expressive enough to fit any data point in the training set [liang2020, ma2018], i.e., the so-called interpolation condition holds, which means that, given $x^{*}\in\arg\min_{x\in\mathbb{R}^{n}}f(x)$, then $x^{*}\in\arg\min_{x\in\mathbb{R}^{n}}f_{i}(x)$ for all $i\in\{1,\dots,N\}$. When the interpolation property holds, SGD methods can be shown to match gradient descent complexity of $\mathcal{O}(\frac{1}{\epsilon})$ in the convex case and its linear rate under strong convexity.
Additionally, the landscape of the training loss of deep models has been studied [liu2022]; while convexity and strong convexity seem not to hold, not even locally, another property appears to hold in large portions of the space: the Polyak-Lojasiewicz (PL) condition [polyak1987introduction]. Very interestingly, the linear convergence rate was shown to hold for SGD even with this weaker property, if interpolation is taken into account [vaswani2019, loizou2021, galli2023].
@@
Loizou et al. [loizou2021] proposed a stochastic variant of the Polyak step size [polyak1987introduction] for guessing the most promising step size for SGD at each iteration. In particular, the Stochastic Polyak step size (SPS) is obtained assuming the optimal function value for the current mini-batch objective $f_{k}$ to be some known value $f_{k}^{*}$ and reachable moving along $d_{k}$; the SPS is consequently obtained, minimizing a first-order model, setting
@@
where $c>0$ is a normalization parameter. In the same work, the bounded variant $\text{SPS}_{\text{max}}$ is also presented, considering $\alpha_{k}=\max\left\{\alpha^{\text{SPS}}_{k},\alpha_{\text{max}}\right\}$ where $\alpha_{\text{max}}>0$ is a suitable upper bound for the step size. The assumption on the value of $f_{k}^{*}$ is reasonable in most cases, as it is linked to the interpolation condition and the non-negative nature of loss functions in most machine learning applications. For instance, it can be set to 0 with unregularized least squares, or logistic losses or by some closed-form rule in the corresponding regularized cases (see, e.g., [loizou2021, Sec. 2.1]).
Inspired by the arguably top performing methods in classical nonlinear optimization, line searches have also been considered to achieve fast convergence in the interpolation regime. The first stochastic line search specifically designed for this scenario was introduced in [vaswani2019]: the extension of an Armijo-type condition [armijo1966] to the stochastic case is shown to lead to the desired linear convergence rate under reasonable assumptions. The line search is structured so that the progress in the current stochastic function is measured, looking for a sufficient decrease that depends on the stochastic gradient direction; formally, the proposed Armijo condition is
@@
In fact, the strict decrease on the current stochastic objective might appear as an unnecessarily strong condition to impose, especially in highly nonlinear and nonconvex problems. This observation was the main motivation, in [galli2023], to introduce a nonmonotone variant of the Armijo type condition. Letting
@@ 3 Integrating Momentum Terms with SLS
It is broadly accepted that the employment of momentum terms in gradient based optimization of finite-sum functions has a considerable positive impact [bottou2018, wright2022optimization]. Specifically, heavy-ball like updates follow rules of the form
@@
where $\beta_{k}$ is a suitable positive scalar. In the stochastic setting, partly repeating the previous update mainly helps at stabilizing the optimization trajectory and at taking larger steps in low-curvature regions. Momentum is thus employed in most state-of-the-art optimizers for large-scale nonconvex finite-sum problems, e.g., it is included in Adam [kingma2015].
@@ 3.1 Existing Strategies and Computational Shortcomings
A generalization of the Stochastic Polyak step size was introduced to determine the appropriate step size when considering a momentum-type direction [wang2023]. In particular, the step size associated with moving average of gradients (MAG) directions of the form $d_{k}=-(1-\beta)\nabla f(x^{k})+\beta(x^{k}-x^{k-1})$, being $\beta\in(0,1)$ a fixed parameter, is defined as
@@
The combination of a momentum-type direction and the use of a stochastic line search was recently discussed [fan2023]. The resulting algorithm, (called MSL_SGDM) revolves around directions of the form (8). In case, for a given value of $\beta$, $d_{k}$ is not a descent direction for the current stochastic function $f_{k}$ (i.e., $d_{k}^{T}\nabla f_{{}_{k}}(x^{k})>0$), it is modified by either setting $\beta=0$ or by reducing $\beta$ until $d_{k}^{T}\nabla f_{k}(x^{k})<0$. Once $d_{k}$ is ensured to be of descent, the step size $\alpha_{k}$ in the update
@@
By means of the same strategy, other families of directions could be considered, as long as $d_{k}$ can be manipulated to satisfy the “stochastic” descent property for the mini-batch; Adam in particular is considered in [fan2023], but Mean Average Gradient (MAG) or Nesterov-type directions could also be handled by this framework in principle.
Although apparently well planned, the strategy from [fan2023] hides a significant drawback, as we will argue hereafter. Surely, it is a given fact that, in the incremental setup, using a momentum-based algorithm somehow allows to mitigate the presence of noise in gradient evaluations caused by the usage of mini-batches.
@@ 3.2 Leveraging Mini-Batch Persistency
In the context of machine learning, this idea translates into using a shared portion of data samples in successive mini-batches. The larger the overlap between mini-batches, the more similar $f_{{k-1}}$ and $f_{k}$ will be. This strategy is referred to as mini-batch persistency, and it was observed to be beneficial even with vanilla SGD algorithms, especially when large batch sizes are considered [fischetti2018].
@@
  • a convex problem - an RBF-kernel classifier to be trained on the ijcnn dataset111available at csie.ntu.edu.tw/ cjlin/libsvmtools/datasets from LIBSVM [libsvm2011];
  • a nonconvex problem - a multi-layer perceptron with 1000 hidden units and ReLU activations to be trained on the MNIST [mnist1998] dataset222available at yann.lecun.com/exdb/mnist/.
@@
A convenient solution to force a 50% overlap in random reshuffling implementations [gurbuzbalaban2021random] of mini-batch GD is to partition training data into $2N_{b}$ disjoint sets $\mathcal{R}^{k}_{1},\dots,\mathcal{R}_{2N_{b}}^{k}$ at each epoch $\kappa$ and work, at iteration $j$ within the epoch, with the subset ${B}_{j}^{\kappa}=\mathcal{R}^{\kappa}_{j-1}\cup\mathcal{R}^{\kappa}_{j}$, so that, for all $j=2,\ldots,2N_{b}$, we have ${B}_{j}^{\kappa}\cap\mathcal{B}_{j+1}^{\kappa}=\mathcal{R}_{j}$ and ${B}_{j}^{\kappa}\cap{B}_{h}^{\kappa}=\emptyset$ for all $h>j+1$. With this approach, each data point is used twice (in subsequent iterations) within an epoch and the number of updates per epoch is doubled. Yet, given the sequential usage of the partitions $\mathcal{R}_{j}$, we can keep data in RAM/VRAM without the need of reading the same data from disk twice per epoch. In other words, when overlap is used, the I/O time from disk required per epoch (which is one of the dominating components in the computational cost of algorithms in large-scale learning) does not increase w.r.t. the non-overlap case.
At this point, we find it useful to show how a 50% mini-batch persistency affects SGD with momentum, Adam [kingma2015], PoNoS [galli2023] and MSL_SGDM [fan2023]. To this aim, we look at the progress of the training loss through time on the MNIST problem already employed above. Each algorithm is run - for 3 different random seeds - for a maximum of 200 epochs and is stopped earlier in case a solution $x^{k}$ s.t. $f(x^{k})<10^{-5}$ is found. Batch sizes of 128 and 512 are tested. The results shown in Figure 2 suggest that a 50% overlap can lead to a significant improvement in the performance of these methods in the tasks considered.
@@ 4 Data-persistent Conjugate Gradient Rules
The conceptual distinction between heavy-ball and CG is subtle and not unambiguously recognized by the optimization community. There is a substantial equivalence in the convex quadratic case, if heavy-ball parameters are chosen optimally [polyak1987introduction]. In the general nonlinear case, CG methods define the search direction by means of a predetermined formula for $\beta_{k}$ and then carry out a Wolfe line search; on the other hand, heavy-ball selects $\alpha_{k}$ and $\beta_{k}$ altogether by some rule. If an Armijo-type line search is employed after choosing $\alpha_{k}$ and $\beta_{k}$ according to some rule, as done, e.g., in [fan2023, lapucci2024globally], we are arguably somewhere in a middle ground. We propose a framework that belongs to this middle ground in the stochastic setting.
@@
In the recent literature, two conjugate-gradient methods for finite-sum problem with convergence guarantees have been proposed [jin2019stochastic, kou2023mini]. The two algorithms are substantially direct descendants of classical CG methods in nonlinear optimization, both possessing convergence properties in expectation. However, they present two major differences w.r.t. what we present here, which actually make their employment very unappealing in deep learning scenarios: they require a Wolfe-type line search, significantly increasing the per-iteration cost, and they employ variance reduction techniques, as the algorithm presented in [jin2019stochastic] requires periodic full batch evaluations of the gradient, and the one presented in [kou2023mini] stores per sample gradients, which have a massive impact on the memory footprint of the method.
@@
Several formulae have been proposed that can be used to compute $\beta_{k}$, that collapse to the same rule when $f$ is quadratic [polyak1987introduction]. Given $y_{k}=\nabla f(x^{k})-\nabla f(x^{k-1})$, some well-known formulations for computing $\beta_{k}$ (cfr. [grippo2023introduction, Sec. 12.5]) are:
@@
The advantage of this strategy is twofold: not only the estimated $\beta_{k}$ will be meaningful when the mini-batch changes, but we also do not need to make additional evaluations, since the four quantities involved in the CG-type formulae would be computed anyway throughout the algorithm. Note that using $\mathcal{R}_{k}$ to compute $\beta_{k+1}$ substantially differs from the strategy used in [jin2019stochastic, kou2023mini], as this allows to compute the CG update rule exactly for the function $f_{\mathcal{R}_{k}}$.

### v4

@@ 1 Introduction
State-of-the-art methods to tackle problems of this form are primarily based on stochastic gradient descent (SGD) [1] and its accelerated variants [2, 3]. The use of stochastic gradient methods allows to leverage data redundancy to approach good solutions while carrying out low-cost iterations. A thorough discussion on stochastic gradient methods and their application in data science is available in [4].
@@
Adaptive SGD approaches [5, 6, 7, 8] have established through the years as solid choices in the considered scenario. In particular, Adam algorithm [8] is now largely considered the preferred optimizer in deep learning settings. In an alternative stream of research, variance-reduced approaches [9, 10, 11] have been studied and shown to possess faster convergence rates than the base algorithm and in line with full-batch gradient descent. However, in practice, these methods require significant computational resources, either needing periodic full gradient computations or high memory usage; for this reason, variance reduction methods are often found to be impractical to use in real-world scenarios with deep architectures and very large datasets. Moreover, the assumptions made to obtain the nice convergence results have been shown not to be satisfied in deep learning scenarios. In fact, no improvement in the convergence rate [12] w.r.t. SGD is obtainable under more realistic assumptions.
@@
In more recent years, the focus turned to the analysis of stochastic gradient methods under a specific assumption that is reasonable with modern deep learning models: the interpolation regime. In this scenario, the convergence rate of SGD was finally shown to match that of full-batch methods [13, 14], giving some justification to the observed behavior of these methods. Moreover, under interpolation assumptions it was possible to design well-grounded strategies to choose the step size in an adaptive manner [15, 16].
@@
Now, one of the most promising developments concerns the use of stochastic line searches. In particular, the classical Armijo rule was adapted to be successfully used in the incremental scenario [17]. More recently, a nonmonotone version of the stochastic Armijo line search was proposed [18]. With these techniques, not only the convergence rate of standard gradient descent is shown to be obtainable, but very encouraging numerical results were also observed.
@@
As widely recognized in the machine learning literature, the addition of momentum terms in the update rule has been known to be beneficial within SGD approaches [19, 20, 21, 22, 23]; typically, this is attributed to the stabilizing effect on the direction and the speed-up in low curvature regions. It is still not clear, however, how this kind of additional term, which modifies the search direction, could be suitably handled in line search based frameworks. A heuristic choice for the step size in presence of momentum terms under interpolation assumptions has been proposed by [24], but no decrease condition is coupled with this strategy. On the other hand, a backtracking strategy has been recently suggested, ensuring the resulting direction is of descent for the considered mini-batch [25]; for this latter approach, however, we will outline massive shortcomings both on the theoretical and on the computational sides.
@@
In this manuscript, we therefore deal with the above challenge. Firstly, we shed light on an intrinsic issue related to the momentum direction in the incremental regime; in order to overcome it, at least partially, we propose to exploit the concept of mini-batch persistency [26], which we also observe being beneficial on its own in certain settings.
@@
We can then focus on a suitable rule to define the parameter $\beta$ associated with the momentum term in the direction update. The proposed method exploits the strong connection of momentum methods and classical nonlinear conjugate gradient algorithms [27, Ch. 12]: in both cases, the search direction is a linear combination of the current gradients and the previous direction, and the border between the two families of approaches is thin. The update rule is combined with safeguarding strategies based on restarting [28] or subspace optimization [29], that allow to guarantee crucial assumptions for the soundness of the resulting algorithm [30].
@@ 2 Preliminaries
The update rule is then given by $x^{k+1}=x^{k}+\alpha_{k}d_{k}$, where the step size $\alpha_{k}$ is often referred to as learning rate and $d_{k}$ is the search direction. A deep presentation of stochastic gradient methods for machine learning tasks is available in [4]. Under regularity assumptions, and using a suitable decreasing sequence of steps $\{\alpha_{k}\}$, the iterative scheme allows to guarantee convergence to stationarity in expectation. The associated worst-case complexity bound is $\mathcal{O}(\frac{1}{\epsilon^{4}})$ in the nonconvex setting [31], which becomes $\mathcal{O}(\frac{1}{\epsilon^{2}})$ and $\mathcal{O}(\frac{1}{\epsilon})$ under convexity and strong convexity assumptions, respectively [4]. These bounds are worse than those obtainable by full-batch methods, which should be thus superior. Yet, the empirical evidence collected training deep machine learning models tells another story.
Recently, some explanation to this gap between theory and numerical observations was found: modern machine learning models are usually expressive enough to fit any data point in the training set [32, 13], i.e., the so-called interpolation condition holds, which means that, given $x^{*}\in\arg\min_{x\in\mathbb{R}^{n}}f(x)$, then $x^{*}\in\arg\min_{x\in\mathbb{R}^{n}}f_{i}(x)$ for all $i\in\{1,\dots,N\}$. When the interpolation property holds, SGD methods can be shown to match gradient descent complexity of $\mathcal{O}(\frac{1}{\epsilon})$ in the convex case and its linear rate under strong convexity.
Additionally, the landscape of the training loss of deep models has been studied [33]; while convexity and strong convexity seem not to hold, not even locally, another property appears to hold in large portions of the space: the Polyak-Lojasiewicz (PL) condition [34]. Very interestingly, the linear convergence rate was shown to hold for SGD even with this weaker property, if interpolation is taken into account [17, 16, 18].
@@
Loizou et al. [16] proposed a stochastic variant of the Polyak step size [34] for guessing the most promising step size for SGD at each iteration. In particular, the Stochastic Polyak step size (SPS) is obtained assuming the optimal function value for the current mini-batch objective $f_{k}$ to be some known value $f_{k}^{*}$ and reachable moving along $d_{k}$; the SPS is consequently obtained, minimizing a first-order model, setting
@@
where $c>0$ is a normalization parameter. In the same work, the bounded variant $\text{SPS}_{\text{max}}$ is also presented, considering $\alpha_{k}=\min\left\{\alpha^{\text{SPS}}_{k},\alpha_{\text{max}}\right\}$ where $\alpha_{\text{max}}>0$ is a suitable upper bound for the step size. The assumption on the value of $f_{k}^{*}$ is reasonable in most cases, as it is linked to the interpolation condition and the non-negative nature of loss functions in most machine learning applications. For instance, it can be set to 0 with unregularized least squares, or logistic losses or by some closed-form rule in the corresponding regularized cases (see, e.g., [16, Sec. 2.1]).
Inspired by the arguably top performing methods in classical nonlinear optimization, line searches have also been considered to achieve fast convergence in the interpolation regime. The first stochastic line search specifically designed for this scenario was introduced in [17]: the extension of an Armijo-type condition [35] to the stochastic case is shown to lead to the desired linear convergence rate under reasonable assumptions. The line search is structured so that the progress in the current stochastic function is measured, looking for a sufficient decrease that depends on the stochastic gradient direction; formally, the proposed Armijo condition is
@@
In fact, the strict decrease on the current stochastic objective might appear as an unnecessarily strong condition to impose, especially in highly nonlinear and nonconvex problems. This observation was the main motivation, in [18], to introduce a nonmonotone variant of the Armijo type condition. Letting
@@ 3 Integrating Momentum Terms with SLS
It is broadly accepted that the employment of momentum terms in gradient based optimization of finite-sum functions has a considerable positive impact [4, 36]. Specifically, heavy-ball like updates follow rules of the form
@@
where $\beta_{k}$ is a suitable positive scalar. In the stochastic setting, partly repeating the previous update mainly helps at stabilizing the optimization trajectory and at taking larger steps in low-curvature regions. Momentum is thus employed in most state-of-the-art optimizers for large-scale nonconvex finite-sum problems, e.g., it is included in Adam [8].
@@ 3.1 Existing Strategies and Computational Shortcomings
A generalization of the Stochastic Polyak step size was introduced to determine the appropriate step size when considering a momentum-type direction [24]. In particular, the step size associated with moving average of gradients (MAG) directions of the form $d_{k}=-(1-\beta)\nabla f(x^{k})+\beta(x^{k}-x^{k-1})$, being $\beta\in(0,1)$ a fixed parameter, is defined as
@@
The combination of a momentum-type direction and the use of a stochastic line search was recently discussed [25]. The resulting algorithm, (called MSL_SGDM) revolves around directions of the form (8). In case, for a given value of $\beta$, $d_{k}$ is not a descent direction for the current stochastic function $f_{k}$ (i.e., $d_{k}^{T}\nabla f_{{}_{k}}(x^{k})>0$), it is modified by either setting $\beta=0$ or by reducing $\beta$ until $d_{k}^{T}\nabla f_{k}(x^{k})<0$. Once $d_{k}$ is ensured to be of descent, the step size $\alpha_{k}$ in the update
@@
By means of the same strategy, other families of directions could be considered, as long as $d_{k}$ can be manipulated to satisfy the “stochastic” descent property for the mini-batch; Adam in particular is considered in [25], but Mean Average Gradient (MAG) or Nesterov-type directions could also be handled by this framework in principle.
Although apparently well planned, the strategy from [25] hides a significant drawback, as we will argue hereafter. Surely, it is a given fact that, in the incremental setup, using a momentum-based algorithm somehow allows to mitigate the presence of noise in gradient evaluations caused by the usage of mini-batches.
@@ 3.2 Leveraging Mini-Batch Persistency
In the context of machine learning, this idea translates into using a shared portion of data samples in successive mini-batches. The larger the overlap between mini-batches, the more similar $f_{{k-1}}$ and $f_{k}$ will be. This strategy is referred to as mini-batch persistency, and it was observed to be beneficial even with vanilla SGD algorithms, especially when large batch sizes are considered [26].
@@
A convenient solution to force a 50% overlap in random reshuffling implementations [39] of mini-batch GD is to partition training data into $2N_{b}$ disjoint sets $\mathcal{R}^{k}_{1},\dots,\mathcal{R}_{2N_{b}}^{k}$ at each epoch $\kappa$ and work, at iteration $j$ within the epoch, with the subset ${B}_{j}^{\kappa}=\mathcal{R}^{\kappa}_{j-1}\cup\mathcal{R}^{\kappa}_{j}$, so that, for all $j=2,\ldots,2N_{b}$, we have ${B}_{j}^{\kappa}\cap\mathcal{B}_{j+1}^{\kappa}=\mathcal{R}_{j}$ and ${B}_{j}^{\kappa}\cap{B}_{h}^{\kappa}=\emptyset$ for all $h>j+1$. With this approach, each data point is used twice (in subsequent iterations) within an epoch and the number of updates per epoch is doubled. Yet, given the sequential usage of the partitions $\mathcal{R}_{j}$, we can keep data in RAM/VRAM without the need of reading the same data from disk twice per epoch. In other words, when overlap is used, the I/O time from disk required per epoch (which is one of the dominating components in the computational cost of algorithms in large-scale learning) does not increase w.r.t. the non-overlap case.
At this point, we find it useful to show how a 50% mini-batch persistency affects SGD with momentum, Adam [8], PoNoS [18] and MSL_SGDM [25]. To this aim, we look at the progress of the training loss through time on the MNIST problem already employed above. Each algorithm is run - for 3 different random seeds - for a maximum of 200 epochs and is stopped earlier in case a solution $x^{k}$ s.t. $f(x^{k})<10^{-5}$ is found. Batch sizes of 128 and 512 are tested. The results shown in Figure 2 suggest that a 50% overlap can lead to a significant improvement in the performance of these methods in the tasks considered.
@@ 4 Data-persistent Conjugate Gradient Rules
The conceptual distinction between heavy-ball and CG is subtle and not unambiguously recognized by the optimization community. There is a substantial equivalence in the convex quadratic case, if heavy-ball parameters are chosen optimally [34]. In the general nonlinear case, CG methods define the search direction by means of a predetermined formula for $\beta_{k}$ and then carry out a Wolfe line search; on the other hand, heavy-ball selects $\alpha_{k}$ and $\beta_{k}$ altogether by some rule. If an Armijo-type line search is employed after choosing $\alpha_{k}$ and $\beta_{k}$ according to some rule, as done, e.g., in [25, 29], we are arguably somewhere in a middle ground. We propose a framework that belongs to this middle ground in the stochastic setting.
@@
In the recent literature, two conjugate-gradient methods for finite-sum problem with convergence guarantees have been proposed [40, 41]. The two algorithms are substantially direct descendants of classical CG methods in nonlinear optimization, both possessing convergence properties in expectation. However, they present two major differences w.r.t. what we present here, which actually make their employment very unappealing in deep learning scenarios: they require a Wolfe-type line search, significantly increasing the per-iteration cost, and they employ variance reduction techniques, as the algorithm presented in [40] requires periodic full batch evaluations of the gradient, and the one presented in [41] stores per sample gradients, which have a massive impact on the memory footprint of the method.
@@
Several formulae have been proposed that can be used to compute $\beta_{k}$, that collapse to the same rule when $f$ is quadratic [34]. Given $y_{k}=\nabla f(x^{k})-\nabla f(x^{k-1})$, some well-known formulations for computing $\beta_{k}$ (cfr. [27, Sec. 12.5]) are:
@@
The advantage of this strategy is twofold: not only the estimated $\beta_{k}$ will be meaningful when the mini-batch changes, but we also do not need to make additional evaluations, since the four quantities involved in the CG-type formulae would be computed anyway throughout the algorithm. Note that using $\mathcal{R}_{k}$ to compute $\beta_{k+1}$ substantially differs from the strategy used in [40, 41], as this allows to compute the CG update rule exactly for the function $f_{\mathcal{R}_{k}}$.